# Tomography for Plasma Imaging: a Unifying Framework for Bayesian Inference

### v2

@@ 1 Introduction
In this paper, we focus on its application to emission tomography, with soft x-ray (SXR) imaging [25] on the TCV tokamak [26] as a specific application. However, the proposed framework is completely general and the discussion readily extends to any diagnostic employing tomography (bolometry [25], visible [27], hard x-ray [28], gamma-ray [29], $\ldots$), on TCV and other machines.
Our main contributions in this paper are:
  • providing a unifying picture of most state of the art plasma emission tomography techniques, highlighting how Bayesian inference principles lie behind most works to date;
  • describing how to obtain, within this general Bayesian framework, estimates of quantities of interest equipped with UQ metrics, by means of well-understood, robust, and efficient inference algorithms;
  • illustrating a concrete application of the entire workflow to SXR imaging on TCV, analyzing the results of a phantom-based study.
@@ 2 Inverse Problems: Bayesian Perspective
where $\mathbf{T}\in\mathbb{R}^{M\times N}$ is a linear forward model describing the acquisition system and $\boldsymbol{\varepsilon}\in\mathbb{R}^{M}$ is the noise corrupting the measurements. Unfortunately, most inverse problems encountered in science are ill-posed [32], i.e., there may be no solution, or a solution may exist but not be unique, or small perturbations of the data $\mathbf{y}$ may greatly affect the solution. In particular, when the problem is under-determined ($N\!>\!M$) and $\mathbf{T}$ is full rank, there are an infinite number of solutions $\mathbf{x}^{*}$ that perfectly satisfy the data $\mathbf{y}$, i.e., such that $\mathbf{T}\mathbf{x}^{*}=\mathbf{y}$. In this Section, we provide a general introduction to regularization-based and Bayesian approaches to inverse problems, highlighting the connections between the two. We refer to [33] for an excellent in-depth discussion of these topics.
@@ 2.2 Bayesian approach
which coincides with the least squares data-fidelity functional from Eq.(3) (up to a multiplicative constant). Thus, such a functional naturally arises from the assumption of additive white Gaussian noise. Furthermore, Laplace-/Gaussian-like priors of the form $\exp\big{(}-\lambda\|\mathbf{G}\mathbf{x}\|_{p}^{p}\big{)}$ with $p\in\{1,2\}$ lead to the regularization terms $R(\mathbf{x})=\|\mathbf{G}\mathbf{x}\|_{p}^{p}$ discussed in Section 2.1. The solution to a penalized least squares problem akin to Eq.(2) can therefore be interpreted as a MAP estimator, where the data-fidelity functional is the negative log-likelihood and the regularization functional is the negative log-prior. This is not the only possible interpretation of penalized likelihood problems, and the choice of $R(\mathbf{x})$ often stems from the empirical observation of its good performance rather than from the belief that $\mathbf{x}$ is actually distributed as the corresponding prior satisfying $-\log p(\mathbf{x})\propto R(\mathbf{x})$ [35]. Nevertheless, the described Bayesian interpretation rationally justifies the formulation of the optimization problem in Eq.(2), providing a sound statistical framework for probabilistic inference.
@@ 3.1 MAP estimation: gradient flow
[…]
@@ 3.2 Posterior sampling: stochastic gradient flow
where $(\mathbf{z}_{k})_{k\geq 0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ are independent realizations of standard $N$-dimen-sional Gaussian vectors. The discrete-time Markov chain in Eq.(14) is the explicit Euler-Maruyama discretization [41] with time step $\Delta t=\tau$ of Eq.(12). This is the stochastic analogous of the explicit Euler method highlighted for Eqs.(10)-(11). For log-posteriors satisfying Eq.(9), the condition $\tau\leq 1/\beta$ guarantees stability of the chain in Eq.(14) and its convergence to a unique stationary distribution $\pi_{\tau}$. Typically, $\pi_{\tau}$ is different from $\pi_{\infty}(\mathbf{x})=p(\mathbf{x}|\mathbf{y})$, but converges to it for $\tau\to 0$ [36]. The bias introduced by such difference can be controlled by reducing $\tau$. In practice, for high-dimensional problems, values close to the upper bound $1/\beta$ are selected, to favor efficiency at the expense of some (typically small) bias [31]. While other variants of ULA have traditionally been preferred [30], recent convergence results [36, 37] have contributed to a resurgence of interest in Eq.(14) for efficient high-dimensional sampling [31].
@@ 4.1 Gradient-penalizing smoothness priors: reaction-diffusion imaging
[…]
@@
[…]
@@
In this work, we always consider diffusion terms such that $R(\mathbf{x})$ is well-defined and given by Eq.(17) for $\mathbf{G}^{T}\mathbf{G}=\boldsymbol{\nabla}^{T}\mathbf{D}\boldsymbol{\nabla}$, i.e., $\mathbf{G}=\sqrt{\mathbf{D}}\boldsymbol{\nabla}$ where by $\sqrt{\mathbf{D}}$ we denote the unique positive semidefinite matrix satisying $\sqrt{\mathbf{D}}\sqrt{\mathbf{D}}=\mathbf{D}$. The operator $\boldsymbol{\nabla}$ is computed through finite differences. Then, as discussed in Sections 2.2 and 3.1, all such regularization terms admit a Bayesian interpretation: we can interpret the steady-state of Eq.(19) as a MAP estimator, with prior $p(\mathbf{x})\propto\exp\big{(}-\lambda R(\mathbf{x})\big{)}$. Moreover, posterior sampling, by means of ULA (Eq.(14)), can be performed. All the plasma tomography-related regularization techniques introduced at the beginning of this Section fit within a Bayesian framework.
@@ 5 Application: SXR Imaging at TCV
Figure 1: SXR diode system at TCV, part of the RADCAM diagnostic. From left to right: phantom model, line of sight (LoS) configuration, and corresponding tomographic data, without ($\mathbf{y}_{true}$) and with ($\mathbf{y}$) noise added.
@@ 5.1 Dataset generation
Figure 2: Phantom generation process: phantoms are obtained summing a randomly generated perturbation $\mathbf{x}_{S}$ to a radially decaying background $\mathbf{b}$. Figs. 2(a)-2(d): initial, intermediate and final anisotropic diffusion iterations performed to compute $\mathbf{x}_{S}$; all iterations are normalized to $1$ for visualization purposes ($\widehat{\mathbf{x}}_{i}=\mathbf{x}_{i}/\max(\mathbf{x}_{i})$). Fig. 2(f): phantom $\mathbf{x}=(\widehat{\mathbf{x}}_{S}+\mathbf{b})/\max(\widehat{\mathbf{x}}_{S}+\mathbf{b})$.
Experimental observations from multiple tokamaks show that SXR emission can exhibit, due to a number of mechanisms, poloidal asymmetries, ring-like structures, peaked profiles, and other features, which can be described as perturbations upon a background SXR emission (see e.g. [10, 3]). SXR tomography algorithms are typically validated on phantoms that feature profiles that are either Gaussian-like, hollow, peaked or exhibit high-field/low-field side (in/out as seen radially) asymmetries [3, 4]. Inspired by these approaches, we construct a large set of $10^{3}$ SXR-like phantoms, each obtained by summing a calculated perturbation to a background profile. In Figure 2, we illustrate the phantom generation process, which we now describe in detail. We consider poloidally symmetric background emissions $\mathbf{b}$, defined as radially decaying flux surface functions whose peak coincides with the magnetic axis. The perturbations are generated by leveraging the diffusion equation analogy from Eq.(19), discussed in Section 4.1. In particular, let $\psi$ be the poloidal magnetic flux function associated to a TCV magnetic equilibrium. Moreover, let $\mathbf{D}=\mathbf{D}(\psi,\alpha)$ be the anisotropic diffusion tensor introduced in Eq.(18), defined as $\mathbf{D}_{n}=\mathbf{e}_{n,\parallel}\mathbf{e}_{n,\parallel}^{T}+\alpha\;\mathbf{e}_{n,\perp}\mathbf{e}_{n,\perp}^{T}$ for pixels $n=1,\ldots,N$; the anisotropic parameter $\alpha<1$ determines the smoothing strength across flux surfaces, i.e., in the direction $\mathbf{e}_{n,\perp}$ oriented in the direction of the local gradient of $\psi$. Then, we obtain a SXR emissivity perturbation $\mathbf{x}_{S}$ by performing $S$ anisotropic diffusion iterations $\mathbf{x}_{n+1}=\mathbf{x}_{n}+\tau\boldsymbol{\nabla}\cdot(\mathbf{D}\boldsymbol{\nabla}\mathbf{x}_{n})$, starting from a profile $\mathbf{x}_{0}$ corresponding to a point-wise source placed at a randomized location within the core. The peak amplitude of the background $\mathbf{b}$ is randomly selected according to a uniform distribution, with bounds defined to satisfy $\max(\mathbf{b})\leq\max(\widehat{\mathbf{x}}_{S})\leq 5\max(\mathbf{b})$, where $\widehat{\mathbf{x}}_{S}=\mathbf{x}_{S}/\max(\mathbf{x}_{S})$.
@@
Finally, the phantom $\mathbf{x}$ is computed as the (normalized to 1) sum of the perturbation $\widehat{\mathbf{x}}_{S}$ and background $\mathbf{b}$, i.e., $\mathbf{x}=(\widehat{\mathbf{x}}_{S}+\mathbf{b})/\max(\widehat{\mathbf{x}}_{S}+\mathbf{b})$. All computations are performed on a $240\times 80$ pixel grid.
To increase the variability in the dataset, we introduce several additional sources of randomness. For each phantom, the magnetic equilibrium is obtained by rescaling and shifting, by random factors, a real TCV equilibrium $\psi_{{}_{TCV}}$. Both the number of anisotropic diffusion iterations and the anisotropic parameter $\alpha$ are randomly selected, with $10^{3}\leq S\leq 10^{4}$, $5\cdot 10^{-4}\leq\alpha\leq 5\cdot 10^{-1}$. In Fig. 3, we show a few of these generated phantoms. Perturbations are less localized for increasing values of $S$, smoother and less sharply anisotropic for larger values of $\alpha$. This large and diverse dataset can be used to test our tomographic pipeline over a wide range of emission profiles. While not fully physically realistic, the class of phantoms defined by the described procedure includes all the phantom types typically used for SXR tomography validation, together with some of the most common experimentally observed emission patterns. This dataset is thus considered suitable for testing tomographic reconstruction algorithms.
@@ 5.2 Phantom Analysis: Results
In Fig.4 and Table 1, we illustrate the results of the Bayesian analysis for three phantoms $\mathbf{x}_{{}_{A}}$, $\mathbf{x}_{{}_{B}}$, $\mathbf{x}_{{}_{C}}$, for tomographic data corrupted according to noise model $N_{2}$. In Figs.4b-4c, we can appreciate how both the MAP and mean estimators successfully reconstruct the main features of the corresponding phantoms; moreover, from Table 1, we observe that $\mathbf{x}_{{}_{MAP}}$ and $\boldsymbol{\mu}_{{}_{ULA}}$ perform similarly well in terms of RMSE and $\delta^{P_{rad}}$. In Figs.4c, we additionally show the true emissivity peak location (red) and its estimated location $\boldsymbol{\mu}_{{}_{ULA}}^{peak}$ (black), together with the rectangular bounding box defined by $\boldsymbol{\mu}_{{}_{ULA}}^{peak}\pm\boldsymbol{\sigma}_{{}_{ULA}}^{peak}$, representing the uncertainty associated to the estimation. The predicted uncertainty $\boldsymbol{\sigma}_{{}_{ULA}}^{peak}$ is found to provide a reliable statistical estimation of the peak location. In Figs.4d, we plot the pixel-wise standard deviation $\boldsymbol{\sigma}_{{}_{ULA}}$. Throughout our phantom analysis, we observe two interesting features: the magnitude of $\boldsymbol{\sigma}_{{}_{ULA}}$ depends on the estimated anisotropic parameter $\alpha$, while its spatial pattern exhibits limited variability across phantoms. Such features can be distinguished in Figs.4d. Indeed, the estimated uncertainty increases when moving from top to bottom of the figure, from the smoother phantom $\mathbf{x}_{{}_{A}}$ to the more anisotropic $\mathbf{x}_{{}_{B}},\mathbf{x}_{{}_{C}}$. Furthermore, for all three phantoms, the standard deviation $\boldsymbol{\sigma}_{{}_{ULA}}$ is higher along two/three specific radial directions. This is particularly apparent in the case of phantom $\mathbf{x}_{{}_{C}}$, due to the plotting scale, but all phantoms show some degree of this behavior. These two observations are ascribed to the uncertainty structure for our close-to-Gaussian reconstruction model in Eq.(24). In the absence of the positivity constraint, $\boldsymbol{\sigma}_{{}_{ULA}}$ would simply be the square root of the diagonal of the posterior covariance matrix $\mathbf{\Sigma}_{post}=(\lambda\boldsymbol{\nabla}^{T}\mathbf{D}(\psi,\alpha)\boldsymbol{\nabla}+\mathbf{T}^{T}\mathbf{I}_{\widehat{\sigma}^{-\!2}}\mathbf{T})^{-1}$, as discussed in Section 4.2: the resulting $\boldsymbol{\sigma}_{{}_{ULA}}$ can then be shown to increase for decreasing values of $\alpha$, and its closed-form expression, for noise models $N_{1}$ and $N_{2}$, does not depend explicitly on the tomographic data $\mathbf{y}$, but only indirectly through the estimated value of $\alpha$. In the case of noise model $N_{3}$, $\boldsymbol{\sigma}_{{}_{ULA}}$ still conserves an explicit dependence on $\mathbf{y}$ through the term $\mathbf{I}_{\widehat{\sigma}^{-\!2}}$, but this dependence mainly affects its magnitude, without significantly affecting its spatial pattern.
@@
$\mathbf{x}_{{}_{A}}$ $2.3\!\cdot\!10^{-2}$ $2.2\!\cdot\!10^{-2}$ $-1.0\%$ $-0.5\%$ $1.3\,cm$ $93.0\%$ $98.4\%$
$\mathbf{x}_{{}_{B}}$ $3.3\!\cdot\!10^{-2}$ $3.3\!\cdot\!10^{-2}$ $+0.1\%$ $+0.4\%$ $4.0\,cm$ $91.6\%$ $98.5\%$
$\mathbf{x}_{{}_{C}}$ $3.9\!\cdot\!10^{-2}$ $4.6\!\cdot\!10^{-2}$ $-0.6\%$ $-0.3\%$ $10.1\,cm$ $87.5\%$ $98.3\%$
@@
$N_{1}$ $3.1\!\cdot\!10^{-2}$ $3.0\!\cdot\!10^{-2}$ $-0.3\%$ $+0.4\%$ $4.3\,cm$ $92.3\%$ $98.5\%$
$N_{2}$ $4.1\!\cdot\!10^{-2}$ $4.2\!\cdot\!10^{-2}$ $-0.5\%$ $+1.0\%$ $5.6\,cm$ $88.8\%$ $98.0\%$
$N_{3}$ $4.1\!\cdot\!10^{-2}$ $4.3\!\cdot\!10^{-2}$ $-1.0\%$ $+0.1\%$ $6.3\,cm$ $90.5\%$ $98.0\%$

### v3

@@ 1 Introduction
In this paper, we focus on the application of Bayesian inference to emission tomography, with soft x-ray (SXR) imaging [25] on the TCV tokamak [26] as a specific application. However, the proposed framework is completely general and the discussion readily extends to any diagnostic employing tomography (bolometry [25], visible [27], hard x-ray [28], gamma-ray [29], $\ldots$), on TCV and other machines.
Our main contributions in this paper are as follows.
  • We provide a unifying picture of most state of the art plasma emission tomography techniques, highlighting how Bayesian inference principles lie behind most works to date.
  • We propose a general Bayesian framework that promotes a novel perspective on plasma tomography, offering a highly adaptable and statistically principled environment which, by design, encompasses previous efforts and can easily be built upon. This framework paves the way for the application to plasma imaging of the many recent advances made by the broader computational imaging community, which include, for example, priors promoting sparsity rather than smoothness and powerful data-driven priors.
  • We describe how to obtain, within the proposed framework, estimates of quantities of interest equipped with uncertainty quantification metrics, by means of well-understood, robust, and efficient inference algorithms. In particular, we propose for plasma tomography the unadjusted Langevin algorithm, a Bayesian inference algorithm that is gaining traction in the imaging community.
  • We illustrate a concrete application of the entire workflow to SXR imaging on TCV, analyzing the results of a phantom-based study.
  • In an effort towards reproducibility and reusability, we provide full open access to the computational routines and data used in this study, making them available to the community as a support for further tomography development and/or validation studies.
@@ 2 Inverse Problems: Bayesian Perspective
where $\mathbf{T}\in\mathbb{R}^{M\times N}$ is a linear forward model describing the acquisition system and $\boldsymbol{\varepsilon}\in\mathbb{R}^{M}$ is the noise corrupting the measurements. Unfortunately, most inverse problems encountered in science are ill-posed [32], i.e., there may be no solution, or a solution may exist but not be unique, or small perturbations of the data $\mathbf{y}$ may greatly affect the solution. In particular, when the problem is under-determined (N>MN\!>\!M) and $\mathbf{T}$ is full rank, there are an infinite number of solutions $\mathbf{x}^{*}$ that perfectly satisfy the data $\mathbf{y}$, i.e., such that $\mathbf{T}\mathbf{x}^{*}=\mathbf{y}$. In this Section, we provide a general introduction to regularization-based and Bayesian approaches to inverse problems, highlighting the connections between the two. We refer to [33] for an excellent in-depth discussion of these topics.
@@ 2.2 Bayesian approach
which coincides with the least squares data-fidelity functional from Eq.(3) (up to a multiplicative constant). Thus, such a functional naturally arises from the assumption of additive white Gaussian noise. Furthermore, Laplace-/Gaussian-like priors of the form $\exp\big(-\lambda\|\mathbf{G}\mathbf{x}\|_{p}^{p}\big)$ with $p\in\{1,2\}$ lead to the regularization terms $R(\mathbf{x})=\|\mathbf{G}\mathbf{x}\|_{p}^{p}$ discussed in Section 2.1. The solution to a penalized least squares problem akin to Eq.(2) can therefore be interpreted as a MAP estimator, where the data-fidelity functional is the negative log-likelihood and the regularization functional is the negative log-prior. This is not the only possible interpretation of penalized likelihood problems, and the choice of $R(\mathbf{x})$ often stems from the empirical observation of its good performance rather than from the belief that $\mathbf{x}$ is actually distributed as the corresponding prior satisfying $\log p(\mathbf{x})\propto R(\mathbf{x})$ [35]. Nevertheless, the described Bayesian interpretation rationally justifies the formulation of the optimization problem in Eq.(2), providing a sound statistical framework for probabilistic inference.
@@ 3.1 MAP estimation: gradient flow
when such a $\beta$ exists, it is either analytically known or can be numerically estimated [^2] For example, in the penalized least squares case $\log\,p(\mathbf{x}|\mathbf{y})=-\frac{1}{2}\|\mathbf{y}-\mathbf{T}\mathbf{x}\|_{2}^{2}-\frac{\lambda}{2}\|\mathbf{G}\mathbf{x}\|_{2}^{2}\;$, it can be shown that it holds $\beta\leq\|\mathbf{T}\|^{2}+\lambda\|\mathbf{G}\|^{2}$, with $\|\cdot\|$ the spectral matrix norm.. Then, we can compute the MAP by running until convergence the iterative algorithm
@@ 3.2 Posterior sampling: stochastic gradient flow
where $(\mathbf{z}_{k})_{k\geq 0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ are independent realizations of standard $N$-dimen-sional Gaussian vectors. The discrete-time Markov chain in Eq.(14) is the explicit Euler-Maruyama discretization [41] with time step $\Delta t=\tau$ of Eq.(12). This is the stochastic analogous of the explicit Euler method highlighted for Eqs.(10)-(11). For log-posteriors satisfying Eq.(9), the condition $\tau\leq 1/\beta$ guarantees stability of the chain in Eq.(14) and its convergence to a unique stationary distribution $\pi_{\infty\,;\,\tau}\,$. Typically, $\pi_{\infty\,;\,\tau}$ is different from $\pi_{\infty}(\mathbf{x})=p(\mathbf{x}|\mathbf{y})$, but converges to it for $\tau\to 0$ [36], i.e., for vanishing discretization error. The bias introduced by such difference can be controlled by reducing $\tau$. In practice, for high-dimensional problems, values close to the upper bound $1/\beta$ are selected, to favor efficiency at the expense of some (typically small) bias [31]. While other variants of ULA have traditionally been preferred [30], recent convergence results [36, 37] have contributed to a resurgence of interest in Eq.(14) for efficient high-dimensional sampling [31].
@@ 4.1 Gradient-penalizing smoothness priors: reaction-diffusion imaging
$\mathbf{D}_{n}=\alpha_{\parallel}\mathbf{e}_{n,\parallel}\mathbf{e}_{n,\parallel}^{T}+\alpha_{\perp}\mathbf{e}_{n,\perp}\mathbf{e}_{n,\perp}^{T}\;\;\mbox{for anisotropic smoothing, with}$
@@
$\mathbf{D}_{n}=W_{nn}(x_{n})\;\mathbf{I}\qquad\qquad\qquad\;\,\mbox{in the isotropic MFI case, with}$
@@
[…]
@@ 5 Application: SXR Imaging at TCV
Figure 1: SXR diode system installed at TCV. From left to right: poloidal cross-section of TCV showing a SXR phantom model $\mathbf{x}$, with magnetic flux surfaces in white; line of sight (LoS) configuration overlaid to phantom $\mathbf{x}$; corresponding tomographic data, without ($\mathbf{y}_{true}$) and with ($\mathbf{y}$) noise added.
@@ 5.1 Dataset generation
Figure 2: Phantom generation process: phantoms are obtained summing a randomly generated perturbation $\mathbf{x}_{S}$ to a radially decaying background $\mathbf{b}$. Figs. 2(a)-2(d): initial, intermediate and final anisotropic diffusion iterations performed to generate the perturbation $\mathbf{x}_{S}$; all iterations are normalized to 11 for visualization purposes ($\widehat{\mathbf{x}}_{i}=\mathbf{x}_{i}/\max(\mathbf{x}_{i})$). Fig. 2(e): radially decaying background $\mathbf{b}$. Fig. 2(f): phantom $\mathbf{x}$, with $\mathbf{x}=(\widehat{\mathbf{x}}_{S}+\mathbf{b})/\max(\widehat{\mathbf{x}}_{S}+\mathbf{b})$.
Experimental observations from multiple tokamaks show that SXR emission can exhibit, due to a number of mechanisms, poloidal asymmetries, ring-like structures, peaked profiles, and other features, which can be described as perturbations upon a background SXR emission (see e.g. [10, 3]). SXR tomography algorithms are typically validated on phantoms that feature profiles that are either Gaussian-like, hollow, peaked or exhibit high-field/low-field side (in/out as seen radially) asymmetries [3, 4]. Inspired by these approaches, we construct a large set of $10^{3}$ SXR-like phantoms, each obtained by summing a calculated perturbation to a background profile. In Figure 2, we illustrate the phantom generation process, which we now describe in detail. We consider poloidally symmetric background emissions $\mathbf{b}$, defined as radially decaying flux surface functions whose peak coincides with the magnetic axis. The perturbations are generated by leveraging the diffusion equation analogy from Eq.(19), discussed in Section 4.1. In particular, let $\psi$ be the poloidal magnetic flux function associated to a TCV magnetic equilibrium. Moreover, let $\mathbf{D}=\mathbf{D}(\psi,\alpha)$ be the anisotropic diffusion tensor introduced in Eq.(18), defined as $\mathbf{D}_{n}=\mathbf{e}_{n,\parallel}\mathbf{e}_{n,\parallel}^{T}+\alpha\;\mathbf{e}_{n,\perp}\mathbf{e}_{n,\perp}^{T}$ for pixels $n=1,\ldots,N$; the anisotropic parameter $\alpha<1$ determines the smoothing strength across flux surfaces, i.e., in the direction $\mathbf{e}_{n,\perp}$ oriented in the direction of the local gradient of $\psi$. Then, we obtain a SXR emissivity perturbation $\mathbf{x}_{S}$ by performing $S$ anisotropic diffusion iterations $\mathbf{x}_{n+1}=\mathbf{x}_{n}+\tau\boldsymbol{\nabla}\cdot(\mathbf{D}\boldsymbol{\nabla}\mathbf{x}_{n})$, starting from a profile $\mathbf{x}_{0}$ corresponding to a point-wise source placed at a randomized location within the core.
@@
Finally, the phantom $\mathbf{x}$ is computed as the (normalized to 1) sum of the normalized perturbation $\widehat{\mathbf{x}}_{S}=\mathbf{x}_{S}/\max(\mathbf{x}_{S})$ and background $\mathbf{b}$, i.e., $\mathbf{x}=(\widehat{\mathbf{x}}_{S}+\mathbf{b})/\max(\widehat{\mathbf{x}}_{S}+\mathbf{b})$. All computations are performed on a $240\times 80$ pixel grid.
To increase the variability in the dataset, we introduce several additional sources of randomness. For each phantom, the magnetic equilibrium is obtained by rescaling and shifting, by random factors, a real TCV equilibrium $\psi_{{}_{TCV}}$. The peak amplitude of the background $\mathbf{b}$ is randomly selected, with bounds defined to satisfy $\max(\mathbf{b})\leq\max(\widehat{\mathbf{x}}_{S})\leq 5\max(\mathbf{b})$. Both the number of anisotropic diffusion iterations and the anisotropic parameter $\alpha$ are randomly selected, with $10^{3}\leq S\leq 10^{4}$, $5\cdot 10^{-4}\leq\alpha\leq 5\cdot 10^{-1}$. Such bounds are picked as adequate, based on some trial runs, to obtain a sufficiently diverse and physically realistic dataset; perturbations are less localized for increasing values of $S$, smoother and less sharply anisotropic for larger values of $\alpha$. In Fig. 3, we show a few of these generated phantoms. This large and diverse dataset can be used to test our tomographic pipeline over a wide range of emission profiles. While not fully physically realistic, the class of phantoms defined by the described procedure includes all the phantom types typically used for SXR tomography validation, together with some of the most common experimentally observed emission patterns [3, 4, 10]. This dataset is thus considered suitable for testing tomographic reconstruction algorithms.
@@ 5.2 Phantom Analysis: Results
In Fig.4 and Table 1, we illustrate the results of the Bayesian analysis for three phantoms $\mathbf{x}_{{}_{A}}$, $\mathbf{x}_{{}_{B}}$, $\mathbf{x}_{{}_{C}}$, for tomographic data corrupted according to noise model $N_{2}$. In Figs.4b-4c, we can appreciate how both the MAP and mean estimators successfully reconstruct the main features of the corresponding phantoms; moreover, from Table 1, we observe that $\mathbf{x}_{{}_{MAP}}$ and $\boldsymbol{\mu}_{{}_{ULA}}$ perform similarly well in terms of RMSE and $\delta^{P_{rad}}$. In Figs.4c, we additionally show the true emissivity peak location (red) and its estimated location $\boldsymbol{\mu}_{{}_{ULA}}^{peak}$ (black), together with the rectangular bounding box defined by $\boldsymbol{\mu}_{{}_{ULA}}^{peak}\pm\boldsymbol{\sigma}_{{}_{ULA}}^{peak}$, representing the uncertainty associated to the estimation. While the estimate $\boldsymbol{\mu}_{{}_{ULA}}^{peak}$ can be less accurate for some specific emissivity patterns, as for the shown phantom $\mathbf{x}_{{{}_{C}}}$, the predicted uncertainty $\boldsymbol{\sigma}_{{}_{ULA}}^{peak}$ is found to consistently provide a reliable statistical estimation of the peak location. In Figs.4d, we plot the pixel-wise standard deviation $\boldsymbol{\sigma}_{{}_{ULA}}$. Throughout our phantom analysis, we observe two interesting features: the magnitude of $\boldsymbol{\sigma}_{{}_{ULA}}$ depends on the estimated anisotropic parameter $\alpha$, while its spatial pattern exhibits limited variability across phantoms. Such features can be distinguished in Figs.4d. Indeed, the estimated uncertainty increases when moving from top to bottom of the figure, from the smoother phantom $\mathbf{x}_{{}_{A}}$ to the more anisotropic $\mathbf{x}_{{}_{B}},\mathbf{x}_{{}_{C}}$. Furthermore, for all three phantoms, the standard deviation $\boldsymbol{\sigma}_{{}_{ULA}}$ is higher along two/three specific radial directions. This is particularly apparent in the case of phantom $\mathbf{x}_{{}_{C}}$, due to the plotting scale, but all phantoms show some degree of this behavior. These two observations are ascribed to the uncertainty structure for our close-to-Gaussian reconstruction model in Eq.(24). In the absence of the positivity constraint, $\boldsymbol{\sigma}_{{}_{ULA}}$ would simply be the square root of the diagonal of the posterior covariance matrix $\mathbf{\Sigma}_{post}=(\lambda\boldsymbol{\nabla}^{T}\mathbf{D}(\psi,\alpha)\boldsymbol{\nabla}+\mathbf{T}^{T}\mathbf{I}_{\widehat{\sigma}^{-\!2}}\mathbf{T})^{-1}$, as discussed in Section 4.2: the resulting $\boldsymbol{\sigma}_{{}_{ULA}}$ can then be shown to increase for decreasing values of $\alpha$, and its closed-form expression, for noise models $N_{1}$ and $N_{2}$, does not depend explicitly on the tomographic data $\mathbf{y}$, but only indirectly through the estimated value of $\alpha$. In the case of noise model $N_{3}$, $\boldsymbol{\sigma}_{{}_{ULA}}$ still conserves an explicit dependence on $\mathbf{y}$ through the term $\mathbf{I}_{\widehat{\sigma}^{-\!2}}$, but this dependence mainly affects its magnitude, without significantly affecting its spatial pattern.
@@
$\mathbf{x}_{{}_{A}}$ $3\!\cdot\!10^{-2} 2.2$$2\!\cdot\!10^{-2}$ $1.0\%$ $0.5\%$ 1.3cm1.3\,cm $93.0\%$ $98.4\%$
$\mathbf{x}_{{}_{B}}$ $3\!\cdot\!10^{-2} 3.3$$3\!\cdot\!10^{-2}$ $0.1\% +$$0.4\%$ 4.0cm4.0\,cm $91.6\%$ $98.5\%$
$\mathbf{x}_{{}_{C}}$ $3.9\!\cdot\!10^{-2}$ $4.6\!\cdot\!10^{-2}$ $0.6\%$ $0.3\%$ 10.1cm10.1\,cm $87.5\%$ $98.3\%$
@@
$N_{1}$ $3.1\!\cdot\!10^{-2}$ $\cdot\!10^{-2} 0.3-0.3\%$ $0.4\%$ 4.3cm4.3\,cm $92.3\%$ $98.5\%$
$N_{2}$ $4.1\!\cdot\!10^{-2}$ $4.2\!\cdot\!10^{-2}$ $0.5\%$ $1.0\%$ 5.6cm5.6\,cm $88.8\%$ $98.0\%$
$N_{3}$ $4.1\!\cdot\!10^{-2}$ $4.3\!\cdot\!10^{-2}$ $\% +0.1+0.1\%$ 6.3cm6.3\,cm $90.5\%$ $98.0\%$